OpenStudy (anonymous):

the sum of two numbers is 17. if one number is subtracted from the other, their difference is -3. i have to write a system of equations with this problem. thanks.

OpenStudy (mathstudent55):

Call one number x and the other y. How do you reporesent the following statement with an equation using x and y? "the sum of two numbers is 17"

OpenStudy (anonymous):

x+y=17

OpenStudy (mathstudent55):

good. You have 1 equation. now we need a second equation.

OpenStudy (anonymous):

x+y=17 x-y=-3

OpenStudy (mathstudent55):

You got it. That's your system of equations.

OpenStudy (anonymous):

ok. I got 7,10. is it right?

OpenStudy (mathstudent55):

x + y = 17 x - y = -3 ------------(add equations) 2x = 14 x = 7 Substitute x = 7 into first eq: 7 + y = 17 y = 10 7 and 10 is the correct answer